# Non-equilibrium charge-vortex duality

### v1

@@ Abstract
Charge vortex dualities in two-dimensional systems provide a fascinating perspective on the symmetries and dynamics that govern the behavior of fields and particles in a planar geometry [1, 2, 3]. This concept fundamentally arises from the theoretical framework of dualities in field theories, which typically express one physical scenario in terms of another, often leading to profound insights about the behavior of systems under varying conditions, such as temperature and coupling constants.
Charge vortex dualities enable transformations between theories describing charged particles and those describing vortices, offering a unique window into understanding phenomena such as superconductivity and the quantum Hall effect. The principle here hinges on representing the interactions or dynamics of one type of excitation the viewpoint of another, fundamentally different type, thus revealing deeper structures of the physical system.
The real power of charge vortex dualities emerges in their ability to simplify complex physical problems. For example, in situations where the behavior of vortices in a system is difficult to compute due to strong coupling effects, the dual theory in terms of charges might be weakly coupled and hence computationally tractable. This inversion of complexity not only aids theoretical calculations but also helps in providing qualitative insights into phase transitions, critical behavior, and topological properties of materials. Such dualities are not just theoretical constructs but have practical implications in designing experiments and interpreting observational data in condensed matter physics, offering a robust toolset for exploring new states of matter in two-dimensional materials.
@@
In this Letter we construct explicit mappings between finite temperature ideal fluids and superfluids, leading to formulations of the effective action for the Euler equations and the Landau-Tisza model as gauge theories. In these theories, vortices are mapped to the charges of gauge theories, broadening our perspective on fluid dynamics and phase transitions in both equilibrium and non-equilibrium frameworks.
Correspondence principle between elasticity and hydrodynamics — As a motivation for the development of dual gauge theory description for fluids we start with the correspondence that emerges between equations of linear elasticity and the Stokes’ equations in hydrodynamics [31, 32].
@@
with the key distinction lying in the nature of the shear modulus: imaginary for fluids and real for solids. We see that the form of elastic and linear fluid equations is exactly the same provided that we substitute elastic coefficients with viscosities. Therefore we conclude that upon the above substitution the dual gauge theory equations in elasticity will describe Stokes flows. Although constructing explicit maps beyond the linear regime is perhaps not possible, we intend to build a systematic procedure for the dual description of fluids and finite-temperature superfluids based on effective actions.
Effective theory of superfluids at finite temperature — We consider a superfluid characterized by its phase field, $\psi$, which signifies the Goldstone mode resulting from $U(1)$ symmetry breaking. Concurrently, we include analogous fields relevant for the description of the normal component to describe interactions within the two-fluid medium. In a setting that is both isotropic and homogeneous, the description of these fields is presented as follows:
@@
Dual formulation — In exploring the effective action description of superfluids, we shift our focus toward developing a dual formulation. This step is motivated by recognizing that boson-vortex type dualities fundamentally rely on an effective action that portrays a Goldstone field, which may exhibit singular behavior. To handle these cases, we utilize canonical variables that impose a set of conservation-like constraints. These constraints are subsequently addressed through the introduction of gauge fields, a strategy that has proven effective in the study of equilibrium superfluids. In these contexts, Goldstone fields arise from the $U(1)$ symmetry and are similarly employed in elastic theories, where they represent various space-time symmetries.
@@
Defects/vortices — The effective field theory we have developed so far overlooks a crucial component of fluid dynamics—vortices. Vortices represent singular configurations in fluid flow, characterized by large-scale rotational movements. These formations inherently break the initial symmetries of the theory by establishing a preferred rest frame for the vortices, which introduces spontaneous symmetry breaking.
@@
Vortex motion on a plane is a classic subject [36, 37]. Motivated by detailed studies of point-vortex dynamics on a plane, it is evident that vortices typically display constrained dynamics. This behavior arises from the conservation of both the dipole moment and the trace of the quadrupole moment [38, 39]. The unconventional symmetry properties of vortices have motivated effective field theory developments in the studies of their macroscopic, many-body properties [40, 41, 42, 43, 44].
@@
[…]
@@
(for a detailed derivation see Appendices). In a superfluid at finite temperature, we observe two distinct species of vortices. One species exists within the superfluid condensate and is described by the $U(1)$ gauge field. The other species corresponds to the normal component of the fluid, which sources the symmetric tensor gauge field. Notably, vortices in the normal component face mobility constraints if the number of vortices is fixed and cannot fluctuate. We note that vortices are not simply related to vorticity field
$\omega=\epsilon^{\alpha\beta\gamma}u_{\alpha}\nabla_{\beta}u_{\gamma}.$ (13)
Physically, the above constraint means that while a single vortex on a plane cannot move, a dipole of vorticity is capable of movement. This observation parallels the symmetry structure seen in point vortex dynamics. However, it’s important to note that in the previous context, the symmetry exists at the level of solutions. In our scenario, the mobility restrictions are inherent to the theory itself. Additionally, we begin with a relativistic effective action, indicating that the symmetry of the state with vortices is analogous to its non-relativistic counterpart. We observe that the singularities within the superfluid component do not exhibit constrained dynamics according to the theory itself. This can be readily seen by looking at the conservation equations
@@
$<A_{y},\psi>=-\frac{iC_{2}}{C_{3}k}\frac{G_{S}\left(1+\frac{\omega^{2}k^{2}}{C_{3}^{2}}\frac{1}{f_{1}(\omega,k)f_{2}(\omega,k)}\right)}{\omega^{2}-(G_{S}/C_{1})k^{2}},$ (16a)
$<A_{y},A_{yy}>=-\frac{\omega k}{C_{3}}\frac{\left(1+\frac{\omega^{2}k^{2}}{C_{3}^{2}}\frac{1}{f_{1}(\omega,k)f_{2}(\omega,k)}\right)}{f_{1}(\omega,k)f_{2}(\omega,k)}.$ (16b)
where we used $f_{1}(\omega,k)=\left(\frac{\omega^{2}}{C_{2}}-\frac{k^{2}}{K_{N}}\right)$ and $f_{2}(\omega,k)=\left(\frac{\omega^{2}}{G_{S}}-\frac{k^{2}}{C_{1}}\right)$ for shortness (for a detailed derivation see Appendices). Finally we note that in a two-fluid mixture there is no static force between vortices in the superfluid component and the normal component. However, there is a dynamical coupling once vortices have moved.
Discussion — We have provided an example of particle-vortex duality that applies to superfluids and fluids at finite temperature. It thus serves a first example of a non-equilibrium duality of the particle-vortex type that maps strong to weak coupling regimes. In consequence our framework offers a unique opportunity to study properties of fluids at strong coupling. Perhaps it can shed light on the classic problems of fluid mechanics, when further genaralized to include viscous effects.
@@ Appendix A APPENDIX A: DUAL ACTION
$\begin{split}&S_{dual}=\frac{1}{2}\int d^{2}xdt\left[G_{S}^{-1}\Sigma_{i}^{2}-K_{S}^{-1}\Sigma_{0}^{2}+D_{ijkl}^{-1}T_{ij}T_{kl}\right.\\ &-K_{N}^{-1}P_{i}^{2}-\frac{2M}{K_{S}G_{N}}T_{ii}\Sigma_{0}+2P_{i}(\partial_{t}\pi_{i})+2\Sigma_{0}(\partial_{t}\pi^{0})\\ &\left.-2\Sigma_{i}(\partial_{i}\pi^{0})-2T_{ij}(\partial_{i}\pi_{j})+\frac{M}{G_{N}}T_{ii}(\partial_{t}\pi^{0})+\frac{2M}{K_{S}}\Sigma_{0}(\partial_{i}\pi_{i})\right],\end{split}$ (19)
@@ Appendix C APPENDIX C: GAUGE TRANSFORMATIONS AND CONTINUITY EQUATIONS
$\begin{split}S_{\text{source}}&=\int d^{2}xdt\left[P_{i}(\partial_{t}\pi^{(s)}_{i})+\Sigma_{0}(\partial_{t}\pi^{0{(s)}})-\Sigma_{i}(\partial_{i}\pi^{0{(s)}})\right.\\ &\left.-T_{ij}(\partial_{i}\pi^{(s)}_{j})+\frac{M}{G_{N}}T_{ii}(\partial_{t}\pi^{0{(s)}})+\frac{M}{K_{S}}\Sigma_{0}(\partial_{i}\pi^{(s)}_{i})\right]\end{split}$
@@
with
$\begin{split}&\rho=\epsilon_{ij}\partial_{i}\partial_{j}\pi^{0(s)}\qquad J_{i}=\epsilon_{ij}(\partial_{j}\partial_{t}-\partial_{t}\partial_{j})\pi^{0(s)}\\ &\tilde{\rho}=\epsilon_{ik}\epsilon_{jl}\partial_{k}\partial_{l}\partial_{i}\pi_{j}^{(s)}\qquad\tilde{J}_{ij}=\epsilon_{ik}\epsilon_{jl}(\partial_{k}\partial_{t}-\partial_{t}\partial_{k})\pi_{l}^{(s)}\\ \end{split}$ (31)
@@ Appendix D APPENDIX D: MOMENTUM SPACE ANALYSIS
[…]
Utilizing the fact that $\tilde{\psi}(-\vec{k},-\omega)=\tilde{\phi}^{*}(\vec{k},\omega)$ and the $\delta$-function identity, we derive the matrix representation of the Lagrangian
@@ Appendix E APPENDIX E: VORTEX DIPOLE - WAVE SOLUTIONS
$\begin{split}&\partial_{i}\mathcal{E}_{i}=G_{S}\rho,\\ &\epsilon_{ij}\partial_{j}\mathcal{B}=C_{1}\left(J_{i}+\frac{1}{G_{S}}\partial_{t}\mathcal{E}_{i}\right)-\frac{C_{1}}{2C_{3}}\epsilon_{ij}\partial_{j}E_{kk},\\ &\partial_{i}^{2}E_{kk}=-C_{2}\tilde{\rho}-\frac{C_{2}}{2C_{3}}\partial_{j}^{2}\mathcal{B},\\ &\epsilon_{ik}\partial_{k}B_{j}=K_{N}\left(\tilde{J}_{ij}+\frac{\delta_{ij}}{C_{2}}E_{kk}\right)+\frac{K_{N}}{2C_{3}}\delta_{ij}\partial_{t}\mathcal{B}.\\ \end{split}$ (37)
@@
$\begin{split}&\partial_{t}(\partial_{i}A_{i})+\partial_{i}^{2}\phi=-G_{s}\rho\\ &\partial_{i}(\partial_{j}A_{j})-\partial_{k}^{2}A_{i}+\frac{C_{1}}{G_{s}}(\partial_{t}^{2}A_{i}+\partial_{t}\partial_{i}\phi)=C_{1}J_{i}+\frac{C_{1}}{2C_{3}}\epsilon_{ij}\partial_{j}\left[\partial_{t}A_{kk}+\partial_{k}\psi_{k}\right]\\ &\partial_{t}(\partial_{j}A_{kk})+\partial^{2}\psi_{j}=C_{2}\tilde{\rho}+\frac{C_{2}}{C_{3}}\epsilon_{kl}\partial_{j}\partial_{k}A_{l}\\ &\partial_{k}\partial_{i}(A_{ki})-\partial_{k}^{2}A_{ii}+\frac{2K_{N}}{C_{2}}(\partial_{t}^{2}A_{kk}+\partial_{t}(\partial_{k}\psi_{k}))=K_{N}\tilde{J}_{ii}+\frac{K_{N}}{C_{3}}\epsilon_{kl}\partial_{k}(\partial_{t}A_{l})\\ \end{split}$ (38)
@@
$\begin{split}&\nabla^{2}\phi-\frac{2C_{1}C_{3}^{2}}{G_{S}(2C_{3}^{2}-C_{1}C_{2})}\partial_{t}^{2}\phi=-G_{S}\rho,\\ &\nabla^{2}A_{i}-\frac{2C_{1}C_{3}^{2}}{G_{S}(2C_{3}^{2}-C_{1}C_{2})}\partial_{t}^{2}A_{i}=-C_{1}\left(J_{i}+\frac{C_{2}}{2C_{3}}\epsilon_{ij}\tilde{\rho}_{j}\right),\\ &\nabla^{2}\psi_{j}-\frac{4K_{N}}{C_{2}}\partial_{t}^{2}\psi_{j}=C_{2}\tilde{\rho}_{j}+\frac{C_{2}}{C_{3}}\epsilon_{kl}\partial_{j}\partial_{k}A_{l},\\ &\nabla^{2}A_{ii}-\frac{4K_{N}}{C_{2}}\partial_{t}^{2}A_{ii}=-2K_{N}\tilde{J}_{ii}-\frac{2K_{N}}{C_{3}}\epsilon_{kl}\partial_{k}(\partial_{t}A_{l}).\\ \end{split}$ (40)

### v2

@@ Abstract
Charge-vortex dualities in two-dimensional systems provide a fascinating perspective on the symmetries and dynamics that govern the behavior of fields and particles in a planar geometry [1, 2, 3]. This concept fundamentally arises from the theoretical framework of dualities in field theories, which typically express one physical scenario in terms of another, often leading to profound insights about the behavior of systems under varying conditions, such as temperature and coupling constants.
Charge-vortex dualities enable transformations between theories describing charged particles and those describing vortices, offering a unique window into understanding phenomena such as superconductivity and the quantum Hall effect. The principle here hinges on representing the interactions or dynamics of one type of excitation the viewpoint of another, fundamentally different type, thus revealing deeper structures of the physical system.
The real power of charge-vortex dualities emerges in their ability to simplify complex physical problems. For example, in situations where the behavior of vortices in a system is difficult to compute due to strong coupling effects, the dual theory in terms of charges might be weakly coupled and hence computationally tractable. This inversion of complexity not only aids theoretical calculations but also helps in providing qualitative insights into phase transitions, critical behavior, and topological properties of materials. Such dualities are not just theoretical constructs but have practical implications in designing experiments and interpreting observational data in condensed matter physics, offering a robust toolset for exploring new states of matter in two-dimensional materials.
@@
In this paper we construct explicit mappings between finite temperature ideal fluids and superfluids, leading to formulations of the effective action for the Euler equations and the Landau-Tisza model as gauge theories. In these theories, vortices are mapped to the charges of gauge theories, broadening our perspective on fluid dynamics and phase transitions in both equilibrium and non-equilibrium frameworks.
I Correspondence principle between elasticity and hydrodynamics
As a motivation for the development of dual gauge theory description for fluids we start with the correspondence that emerges between equations of linear elasticity and the Stokes’ equations in hydrodynamics [31, 32].
@@ I Correspondence principle between elasticity and hydrodynamics
with the key distinction lying in the nature of the shear modulus: imaginary for fluids and real for solids. We see that the form of elastic and linear fluid equations is exactly the same provided that we substitute elastic coefficients with viscosities. Therefore we conclude that upon the above substitution the dual gauge theory equations in elasticity will describe Stokes flows. Although constructing explicit maps beyond the linear regime is more challenging (see [33] for an example of a non-linear map), we intend to build a systematic procedure for the dual description of fluids and finite-temperature superfluids based on effective actions.
II Effective theory of superfluids at finite temperature
We consider a superfluid characterized by its phase field, $\psi$, which signifies the Goldstone mode resulting from $U(1)$ symmetry breaking. Concurrently, we include analogous fields relevant for the description of the normal component to describe interactions within the two-fluid medium. In a setting that is both isotropic and homogeneous, the description of these fields is presented as follows:
@@ II Effective theory of superfluids at finite temperature
III Dual formulation
In exploring the effective action description of superfluids, we shift our focus toward developing a dual formulation. This step is motivated by recognizing that boson-vortex type dualities fundamentally rely on an effective action that portrays a Goldstone field, which may exhibit singular behavior. To handle these cases, we utilize canonical variables that impose a set of conservation-like constraints. These constraints are subsequently addressed through the introduction of gauge fields, a strategy that has proven effective in the study of equilibrium superfluids. In these contexts, Goldstone fields arise from the $U(1)$ symmetry and are similarly employed in elastic theories, where they represent various space-time symmetries.
@@ III Dual formulation
IV Defects/vortices
The effective field theory we have developed so far overlooks a crucial component of fluid dynamics—vortices. Vortices represent singular configurations in fluid flow, characterized by large-scale rotational movements. These formations inherently break the initial symmetries of the theory by establishing a preferred rest frame for the vortices, which introduces spontaneous symmetry breaking.
@@ IV Defects/vortices
Vortex motion on a plane is a classic subject [37, 38]. Motivated by detailed studies of point-vortex dynamics on a plane, it is evident that vortices typically display constrained dynamics. This behavior arises from the conservation of both the dipole moment and the trace of the quadrupole moment [39, 40]. The unconventional symmetry properties of vortices have motivated effective field theory developments in the studies of their macroscopic, many-body properties [41, 42, 43, 44, 45, 33].
@@
$S_{\text{defect}}=\int d^{2}xdt\left(-\rho\phi+A_{i}J_{i}-\tilde{\rho}\psi+A_{ij}\tilde{J}_{ij}\right),$ (10)
@@
(for a detailed derivation see Appendices). In a superfluid at finite temperature, we observe two distinct species of vortices. One species exists within the superfluid condensate and is described by the $U(1)$ gauge field. The other species corresponds to the normal component of the fluid, which sources the symmetric tensor gauge field. Notably, vortices in the normal component face mobility constraints if the number of vortices is fixed and cannot fluctuate. [^1] We note that vortices are not simply related to vorticity field $\omega=\epsilon^{\alpha\beta\gamma}u_{\alpha}\nabla_{\beta}u_{\gamma}.$ Physically, the above constraint means that while a single vortex on a plane cannot move, a dipole of vorticity is capable of movement. This observation parallels the symmetry structure seen in point vortex dynamics. However, it’s important to note that in the previous context, the symmetry exists at the level of solutions. In our scenario, the mobility restrictions are inherent to the theory itself. Additionally, we begin with a relativistic effective action, indicating that the symmetry of the state with vortices is analogous to its non-relativistic counterpart. We observe that the singularities within the superfluid component do not exhibit constrained dynamics according to the theory itself. This can be readily seen by looking at the conservation equations
@@
$A_{y},\psi>=-\frac{iC_{2}}{C_{3}k}\frac{G_{S}\left(1+\frac{\omega^{2}k^{2}}{C_{3}^{2}}\frac{1}{f_{1}(\omega,k)f_{2}(\omega,k)}\right)}{\omega^{2}-(G_{S}/C_{1})k^{2}},$ (15a)
$A_{y},A_{yy}>=-\frac{\omega k}{C_{3}}\frac{\left(1+\frac{\omega^{2}k^{2}}{C_{3}^{2}}\frac{1}{f_{1}(\omega,k)f_{2}(\omega,k)}\right)}{f_{1}(\omega,k)f_{2}(\omega,k)}.$ (15b)
where we used $f_{1}(\omega,k)=\left(\frac{\omega^{2}}{C_{2}}-\frac{k^{2}}{K_{N}}\right)$ and $f_{2}(\omega,k)=\left(\frac{\omega^{2}}{G_{S}}-\frac{k^{2}}{C_{1}}\right)$ for shortness (for a detailed derivation see Appendices). Finally we note that in a two-fluid mixture there is no static force between vortices in the superfluid component and the normal component. However, there is a dynamical coupling once vortices have moved. Specifically, the mixed correlators $\langle A_{y},\psi\rangle$ and $\langle A_{y},A_{yy}\rangle$ indicate that fluctuations in the superfluid $U(1)$ gauge field $A_{y}$ are not dynamically independent of the tensor gauge fields associated with the normal component. This interdependence arises due to the coupling term $M$ in the original Lagrangian, and manifests through the $C_{3}$-dependent contributions in the correlation functions.
Physically, these mixed correlators imply that a moving vortex in one component (e.g., the superfluid) can induce gauge field responses in the other component (e.g., the normal fluid). Although the two-fluid system exhibits no static interaction between vortices residing in different components, the non-vanishing of these mixed two-point functions highlights a non-trivial dynamical coupling.
V Discussion
We have provided an example of particle-vortex duality that applies to superfluids and fluids at finite temperature. It thus serves a first example of a non-equilibrium duality of the particle-vortex type that maps strong to weak coupling regimes. In consequence our framework offers a unique opportunity to study properties of fluids at strong coupling. Perhaps it can shed light on the classic problems of fluid mechanics, when further genaralized to include viscous effects.
@@ Appendix A APPENDIX A: DUAL ACTION
$\begin{split}&S_{\text{dual}}=\frac{1}{2}\int d^{2}xdt\left[G_{S}^{-1}\Sigma_{i}^{2}-K_{S}^{-1}\Sigma_{0}^{2}+D_{ijkl}^{-1}T_{ij}T_{kl}\right.\\ &-K_{N}^{-1}P_{i}^{2}-\frac{2M}{K_{S}G_{N}}T_{ii}\Sigma_{0}+2P_{i}(\partial_{t}\pi_{i})+2\Sigma_{0}(\partial_{t}\pi^{0})\\ &\left.-2\Sigma_{i}(\partial_{i}\pi^{0})-2T_{ij}(\partial_{i}\pi_{j})+\frac{M}{G_{N}}T_{ii}(\partial_{t}\pi^{0})+\frac{2M}{K_{S}}\Sigma_{0}(\partial_{i}\pi_{i})\right],\end{split}$ (18)
@@ Appendix C APPENDIX C: GAUGE TRANSFORMATIONS AND CONTINUITY EQUATIONS
[…]
@@
$S_{\text{source}}=\int d^{2}xdt\left(-\rho\phi+A_{i}J_{i}-\tilde{\rho}\psi+A_{ij}\tilde{J}_{ij}\right)$
with
@@ Appendix D APPENDIX D: MOMENTUM SPACE ANALYSIS
$\begin{split}&\phi=\frac{1}{(2\pi)^{3}}\int d^{2}kd\omega e^{i(\vec{k}\vec{r}-\omega t)}\tilde{\phi}(\vec{k},\omega),\qquad\psi=\frac{1}{(2\pi)^{3}}\int d^{2}kd\omega e^{i(\vec{k}\vec{r}-\omega t)}\tilde{\psi}(\vec{k},\omega),\\ A_{i}&=\frac{1}{(2\pi)^{3}}\int d^{2}kd\omega e^{i(\vec{k}\vec{r}-\omega t)}\tilde{A}_{i}(\vec{k},\omega),\qquad A_{ij}=\frac{1}{(2\pi)^{3}}\int d^{2}kd\omega e^{i(\vec{k}\vec{r}-\omega t)}\tilde{A}_{ij}(\vec{k},\omega).\end{split}$ (32)
Utilizing the fact that $\tilde{\phi}(-\vec{k},-\omega)=\tilde{\phi}^{*}(\vec{k},\omega)$ and the $\delta$-function identity, we derive the matrix representation of the Lagrangian
@@ Appendix E APPENDIX E: VORTEX DIPOLE - WAVE SOLUTIONS
[…]
@@
[…]
@@
[…]